Find 𝑑𝑦 𝑑π‘₯ if 𝑦 = 𝑒 3 βˆ’ 3𝑒 2 + 1 and 𝑒 = π‘₯ 2 + 2Β 


1
Expert's answer
2022-03-30T08:13:04-0400

"\frac{dy}{dx}=\frac{dy}{du}\frac{du}{dx}=\left( 3u^2-6u \right) \cdot 2x=6x\left( \left( x^2+2 \right) ^2-2\left( x^2+2 \right) \right) =\\=6x\left( x^4+2x^2 \right)"